OpenStudy (anonymous):

given a triangle with b=4, c=10, and A=84 degrees, what is the length of a? a. 14.1 b. 10.4 c. 12.5 d. 13.7

OpenStudy (jdoe0001):

hmmm have you cover the Law of Cosines yet?

OpenStudy (anonymous):

Yes @jdoe0001

OpenStudy (jdoe0001):

then is pretty much just a matter of plug and chug =) \(\bf \textit{Law of Cosines}=a^2=b^+c^2-2bc\cdot cos(A)\\ \quad \\ c=\sqrt{a^2=b^+c^2-2bc\cdot cos(A)}\)

OpenStudy (jdoe0001):

wops, darn. so much for my paste =)

OpenStudy (jdoe0001):

\(\bf \textit{Law of Cosines}=a^2=b^2+c^2-2bc\cdot cos(A)\\ \quad \\ a=\sqrt{b^2+c^2-2bc\cdot cos(A)}\)

OpenStudy (anonymous):

i dont know what a is tho..

OpenStudy (jdoe0001):

yes, but you do you know the other factors, "b" and "c" and angle A

OpenStudy (anonymous):

ya ok. let me figure it out and ill come back to ya

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

i got b @jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\bf a=\sqrt{4^2+10^2-2(4)(10)\cdot cos(84^o)} \approx 10.374860140\)